# 22footnotetext: Corresponding author

### v1

@@ 3.5.1. Medical Image Segmentation
Khan et al. (Khan et al., 2024) proposed a convolution and attention-free segmentation mamba based model named CAMS-Net. The authors proposed NC-Mamba block, which differs from normal Mamba block by not using $1D$ convolution. Instead, representations are directly projected into SSM with SiLU activation function. The linearly interconnected factorized mamba block manipulates the input representation using a linear layer with SiLU. When using LIFM block, representations of height and width are combined into a single dimension and later reshaped to their original form. Therefore, LIFM block operates on two levels: Mamba Channel Aggregator (MCA), which works on channel level, and Mamba Spatial Aggregator (MSA), which operates on the combined height and width dimension. CS-IF blocks employs a combination of MCA and MSA blocks in each unit. Forward and backward scans are subsequently applied with shared weights. The initial part of segmentation network includes an MCA block combined with sinusoidal position embedding. Downsampling in encoder is performed using $2\times 2$ average pooling. CS-IF blocks are employed in bottleneck and last encoder part. The decoder also uses MCA blocks, with upsampling performed by $2\times 2$ bi-linear transformation. Overall, this paper aimed to present a convolution and attention-free methodology rather than achieving state-of-the-art results. However, CAMS-Net demonstrates superior or second-best performance compared to convolution and transformer-based methods on CMR×Recon MICCAI-2023 dataset.
Yuan et al. (Yuan et al., 2024) presented U-Net-based segmentation network tailored for skin lesion segmentation. This architecture differs from other mamba based U-Nets by omitting SAB and CAB blocks in skip connections. It features six encoder layers followed by downsampling with convolution layers, and six decoder layers with upsampling through convolution. The initial and final layers of both encoder and decoder utilize convolution layers. Aside from these initial and final layers, MUCM-Net integrates UCM-Net blocks and Mamba blocks by adding their representations together. There are two MUCM-Net configurations:
  1. (1)
    MUCM-Net(1-patch pipeline): This version has a single Mamba block where the feature passes directly through the Mamba block and is then added to the UCM-Net block.
  2. (2)
    Another configuration splits channels into N parts, each processed by separate Mamba blocks, which are then concatenated and passed to the UCM-Net block, with N being 2, 4, or 8.
MUCM-Net was evaluated on ISIC-2017 and ISIC-2018 datasets, demonstrating performance that matched or exceeded that of UCM-Net and other mamba-based architectures such as LightM-UNet by Liao et al. (Liao et al., 2024) and VM-UNet by Ruan et al. (Ruan and Xiang, 2024).
3.5.2. Medical Image Classification
Figure 23. Work Flow of Classification Task
Classification in medical imaging refers to categorizing images into different classes, such as distinguishing between benign and malignant lesions, or identifying different types of diseases. Figure 23 shows the workflow of Mamba-based models in medical image classification task. Table 4 summarizes classification models and their parameters, along with descriptions and information regarding code availability. Some mamba-based architectures applied in this domain include:
@@ 3.5.2. Medical Image Classification
Gong et al. (Gong et al., 2024) introduced nnMamba method for medical image classification that integrates strengths of CNNs and SSMs. Traditional CNNs lack in capturing long-range dependencies due to their local receptive fields, while transformers, though capable of modeling global context, are computationally intensive for 3D medical images. nnMamba addresses these challenges by incorporating Mamba-In-Convolution with Channel-Spatial Siamese (MICCSS) block, which effectively combines long-range dependency modeling and local feature extraction . nnMamba introduces channel-scaling and channel-sequential learning. Channel-scaling adjusts the importance of different feature channels, while channel-sequential learning processes features in a sequential manner to capture complex dependencies. nnMamba’s backbone architecture leverages MICCSS block to maintain computational efficiency while achieving superior performance in 3D medical imaging tasks.
Table 4. Overview of Classification Models
@@
nnMamba (Gong et al., 2024) 15.55 Res-Mamba Alzheimer prediction and Landmark detection
MedMamba (Yue and Li, 2024) 15.2 VMamba Multiple disease classification
Vim4Path (Nasiri-Sarvi et al., 2024) 7 ViM Breast cancer prediction in WSI
Microscopic-Mamba (Zou et al., 2024) 1.59 PEVM Microscopic image classification
Yue et al. (Yue and Li, 2024) propose MedMamba, a novel method for medical image classification that leverages modern SSM, particularly inspired by VMamba. MedMamba addresses limitations of CNN and transformers by combining local feature extraction capability of convolutional layers with long-range dependency modeling of SSMs, thereby achieving efficient and effective classification of medical images. The architecture of MedMamba includes a patch embedding layer, SS-Conv-SSM blocks, and patch merging layers. Figure 24 shows the architecture of MedMamba where it incorporates SS-Conv-SSM blocks. These blocks use a dual-branch approach to separately process and merge features from convolutional and SSM pathways, incorporating a 2D-selective scan (SS2D) for comprehensive feature extraction. Extensive experimental analysis on multiple datasets demonstrate that MedMamba surpasses existing methods in terms of accuracy, establishing a new baseline for medical image classification. The study demonstrates the potential of SSM-based models in medical applications along with the outline of future research directions which includes optimization and integration of explainable AI.
Figure 24. Architecture of MedMamba (Yue and Li, 2024)
Nasiri et al. (Nasiri-Sarvi et al., 2024) used Vision Mamba (ViM) in its architecture for learning representations in histopathology images. Vim4path performs self supervised learning (SSL) within DINO framework (Caron et al., 2021). It extracts image patches from Whole Slide Images (WSIs) and trains ViM encoder using DINO without the availability of labels. ViM is compared with ViT in slide-level and patch-level classification tasks in Camelyon16 dataset. It demonstrates superior performance of ViM at smaller scales and competitive performance at larger scales. The analysis shows that ViM mimics the pathologist’s diagnostic workflow better than ViTs. The paper highlights potential of ViM in practical diagnostic applications and its effectiveness in less computation scenarios which establishes it as a promising tool for computational pathology.
Microscopic-Mamba proposed by Zou et al. (Zou et al., 2024) consists of following components in sequence: patch embedding, four blocks of Hybrid-Conv-SSM each followed by a patch merging block, Global pooling, $1\times 1$ convolution (PW convolution) and finally a fully connected layer for classification. Hybrid-Conv-SSM block splits the representation channel wise which is then passed through two different batches. First branch is the Conv branch which contains depth wise and point wise convolution. Second branch is SSM branch which contains Parallel Efficient Vision Mamba (PEVM) block. Microscopic-Mamba has three variants in terms of parameters such as tiny, small, base. Tiny has 4.32M parameters, small has 4.97M parameters and base has 8.37M parameters. Microscopic-Mamba achieves state-of-the-art results on five different classification datasets outperforming transformer-based methods, mamba-based methods and other hybrid approaches.
3.5.3. Medical Image Restoration/ Reconstruction
Figure 25. Work Flow of Restoration/Reconstruction Task
@@ 3.5.3. Medical Image Restoration/ Reconstruction
Restoration is an application in medical imaging which is used to improve the quality of images that may be corrupted or distorted due to factors such as noise, low resolution and blurring. Reconstruction is a mathematical process that converts raw medical data into target image. Table 5 provides an overview of Mamba based models applied in medical image restoration/reconstruction along with parameters, descriptions and the availability of code. Figure 25 shows the workflow of Mamba-based models in medical image restoration/reconstruction task. Key Mamba-based architectures used in this application are explained as follows.
Table 5. Overview of Restoration Models
@@
MambaMIR (Huang et al., 2024b) - Mamba MRI-CT image reconstruction
FDVM-Net (Zheng and Zhang, 2024) - Mamba Endoscopic image reconstruction
MambaDFuse (Li et al., 2024b) - Mamba Fused reconstruction with MRI,CT,PET images
FusionMamba (Xie et al., 2024a) - DVSS Dynamic Feature Enhancement and Image fusion
@@
Zheng et al. (Zheng and Zhang, 2024) presents FDVM-Net which is designed for endoscopic exposure correction and it leverages frequency-domain reconstruction to achieve high-quality image restoration. To capture spatial features and global dependencies, FDVM-Net combines Mamba and convolutional blocks. This combination serves as the foundation for a dual-path network architecture, which separates processing of the image’s phase and amplitude information. Furthermore, FDVM-Net applies a frequency domain cross-attention module to improve the performance of network.
Huang et al. (Huang et al., 2024b) introduced MambaMIR for medical image reconstruction and uncertainty estimation. Figure 26 shows MambaMIR which incorporates an Arbitrary-Masked State Space (AMSS) block with Monte Carlo dropout. AMSS is inspired by Mamba which comprises of AMS6 block, a gating linear layer, depth-wise convolution layer, and SiLU activation function. Residual connections used in MambaMIR ensures efficient and stable training. This helps MambaMIR to achieve robust performance in medical image reconstruction tasks while providing reliability in reconstruction using uncertainty estimation.
Figure 26. Architecture of MambaMIR (Huang et al., 2024b)
@@
MambaDFuse introduced by Li et al. (Li et al., 2024b) incorporates Multimodal Mamba blocks which effectively merges features from different modality into a unified representation. This helps to obtain an informative fused image as output. MambaDFuse is a dual-phase model that integrates complementary information from diverse imaging modalities into a single image. MambaDFuse utilizes the combined strengths of convolutional layers and Mamba blocks to capture a spectrum of features. It ranges from low-level details to high-level semantic information.
@@
Xie et al. (Xie et al., 2024a) proposed FusionMamba to address limitations of channel redundancy and limited local enhancement capabilities in existing image fusion methods. FusionMamba uses dynamic convolution and channel attention mechanisms to enhance the model’s ability to capture global context and local features within the images. FusionMamba leverages strengths of Mamba, dynamic feature enhancement and cross-modal fusion techniques. FusionMamba explores internal features and relationships between different image modalities. This results in dynamically enhanced texture details, better differences between modalities and improved ability to capture correlations while simultaneously reducing redundant information.
3.5.4. Medical Image Registration
@@
Image registration is the process involves aligning two or more images of the same scene captured at different times, from different viewpoints, or by different sensors. In medical image registration, this typically involves aligning a fixed volume, such as an intraoperative CT scan with a moving volume, such as a preoperative MRI to accurately overlay anatomical structures from different imaging modalities. This step is crucial for diagnosis, treatment planning, and monitoring in medical applications. There are two main types of image registration based on transformation models: linear registration and deformable (non-rigid) registration. Linear registration includes methods such as rigid and affine registration. Rigid registration aligns images using only translations and rotations. It assumes that shapes and sizes of objects in images remain constant. Affine registration includes translation, rotation, scaling, and shearing. This method can handle uniform changes in scale and non-orthogonal angles. Deformable registration refers to the alignment of images using transformation models that allow for local deformations.
Figure 27. Work Flow of Registration Task
Deformable registration can handle complicated transformations which makes it appropriate for aligning MR and CT scans when the patient’s anatomy may have changed between scans. Table 6 describes registration models and their parameters along with information on their descriptions and code availability. Figure 27 shows the workflow of Mamba-based models in medical image registration. Some mamba-based models relevant to this area include:
Table 6. Overview of Registration Models
@@
MambaMorph (Guo et al., 2024) 7.59 Mamba 3D MRI-CT registration
VMambaMorph (Wang et al., 2024c) 9.64 VMamba 3D MRI-CT registration
@@
Guo et al. (Guo et al., 2024) proposed a multi-modality deformable registration framework designed to process the medical MR-CT deformable registration dataset. Inspired by Chen et al. (Chen et al., 2022), which utilizes Swin Transformers in encoder of registration module, MambaMorph instead uses Mamba blocks to capture long-range spatial relationships while optimizing memory utilization. Figure 28 shows that before registration module, a simple fine-grained feature extractor (U-Net with one downsampling step) is utilized to maximize retention of local information.
Figure 28. Architecture of MambaMorph (Guo et al., 2024)
@@
Wang et al. (Wang et al., 2024c) transformed the 2D VSS of VMamba into a 3D volumetric feature processing framework. This hybrid architecture combined VMamba with CNN (U-Net), to accurately estimate the deformation field. To address the challenges posed by complex motion and diverse structures in image registration, VMambaMorph introduces a recursive registration framework. Furthermore, it employs a weight-sharing fine-grained feature extractor to extract features across divergent volumes. MambaMorph’s inadequacy in fully exploiting visual features from complex motion and diverse structures of images or volumes has been surpassed by VMambaMorph.
3.5.5. Miscellaneous

### v3

@@ 3.5.1 Medical Image Segmentation
Khan et al. [89] proposed a convolution and attention-free segmentation mamba based model named CAMS-Net. The authors proposed NC-Mamba block, which differs from normal Mamba block by not using $1D$ convolution. Instead, so the representations are directly projected into SSM with SiLU activation function. When using LIFM block, representations of height and width are combined into a single dimension and later reshaped to their original form. Therefore, LIFM block operates on two levels: Mamba Channel Aggregator (MCA), which works on channel level, and Mamba Spatial Aggregator (MSA), which operates on the combined height and width dimension. The initial part of segmentation network includes an MCA block combined with sinusoidal position embedding. Downsampling in encoder is performed using $2\times 2$ average pooling. CS-IF blocks are employed in bottleneck and last encoder part. The decoder also uses MCA blocks, with upsampling performed by $2\times 2$ bi-linear transformation. Yuan et al. [90] presented U-Net-based segmentation network tailored for skin lesion segmentation. This architecture differs from other mamba based U-Nets by omitting SAB and CAB blocks in skip connections. It features six encoder layers followed by downsampling with convolution layers, and six decoder layers with upsampling through convolution. The initial and final layers of both encoder and decoder utilize convolution layers. Aside from these initial and final layers, MUCM-Net integrates UCM-Net blocks and Mamba blocks by adding their representations together.
Section 3.5.1 focuses on Mamba-based models such as LightM-UNet [80] and Mamba-UNet [78]. Comparing with Transformer-based models such as Swin UNETR [156] and Swin-UNet [131]. LightM-UNet incorporates a Vision Mamba (RVM) layer with residual connections to reduce model complexity, while Swin UNETR relies on Swin Transformer blocks. On the LiTS [55] dataset, LightM-UNet used 1.87M parameters and 457.62 GFLOPs, whereas Swin UNETR required 61.99M parameters and 1570.32 GFLOPs. Similarly, on the Montgomery & Shenzhen [56] datasets, LightM-UNet used just 1.09M parameters and 267.19 GFLOPs, compared to Swin UNETR’s 25.12M parameters and 909.26 GFLOPs. This shows that LightM-UNet is more efficient even with the low computation without reducing the performance. For multi-organ segmentation, Mamba-UNet was evaluated on the ACDC [54] and Synapse CT datasets. While comparing Mamba-UNet with Swin-UNet,the results shows improved performance: on ACDC, Mamba-UNet achieved a Dice score of 0.9281 and Hausdorff Distance (HD) of 2.464, compared to Swin-UNet’s 0.9188 and 3.1817, respectively. On Synapse CT, Mamba-UNet scored 0.6429 Dice and 24.47 HD, outperforming Swin-UNet’s 0.6178 and 30.54. These outcomes highlight the ability of Mamba-based architectures to provide better accuracy with a simple architecture.
3.5.2 Medical Image Classification
Figure 17: Work Flow of Classification Task
@@ 3.5.2 Medical Image Classification
Classification in medical imaging refers to categorizing images into different classes, such as distinguishing between benign and malignant lesions, or identifying different types of diseases. Figure 17 shows the workflow of Mamba-based models in medical image classification task. Table 4 summarizes classification models and their parameters, along with descriptions and information regarding code availability. Some mamba-based architectures applied in this domain include: Gong et al. [20] introduced nnMamba method for medical image classification that integrates strengths of CNNs and SSMs. Traditional CNNs lack in capturing long-range dependencies due to their local receptive fields, while transformers, though capable of modeling global context, are computationally intensive for 3D medical images. nnMamba addresses these challenges by incorporating Mamba-In-Convolution with Channel-Spatial Siamese (MICCSS) block, which effectively combines long-range dependency modeling and local feature extraction . nnMamba introduces channel-scaling and channel-sequential learning. Channel-scaling adjusts the importance of different feature channels, while channel-sequential learning processes features in a sequential manner to capture complex dependencies. nnMamba’s backbone architecture leverages MICCSS block to maintain computational efficiency while achieving superior performance in 3D medical imaging tasks.
Table 4: Overview of Classification Models
@@
nnMamba [20] 15.55 Res-Mamba Alzheimer prediction and Landmark detection
MedMamba [21] 15.2 VMamba Multiple disease classification
Vim4Path [22] 7 ViM Breast cancer prediction in WSI
Microscopic-Mamba [77] 1.59 PEVM Microscopic image classification
Yue et al. [21] propose MedMamba, a novel method for medical image classification that leverages modern SSM, particularly inspired by VMamba. The architecture of MedMamba includes a patch embedding layer, SS-Conv-SSM blocks, and patch merging layers. These blocks use a dual-branch approach to separately process and merge features from convolutional and SSM pathways, incorporating a 2D-selective scan (SS2D) for comprehensive feature extraction. The study demonstrates the potential of SSM-based models in medical applications along with the outline of future research directions which includes optimization and integration of explainable AI. Nasiri et al. [22] used Vision Mamba (ViM) in its architecture for learning representations in histopathology images. Vim4path performs self supervised learning (SSL) within DINO framework [141]. It extracts image patches from Whole Slide Images (WSIs) and trains ViM encoder using DINO without the availability of labels. The paper highlights potential of ViM in practical diagnostic applications and its effectiveness in less computation scenarios which establishes it as a promising tool for computational pathology. Microscopic-Mamba proposed by Zou et al. [77] consists of following components in sequence: patch embedding, four blocks of Hybrid-Conv-SSM each followed by a patch merging block, Global pooling, $1\times 1$ convolution (PW convolution) and finally a fully connected layer for classification. Hybrid-Conv-SSM block splits the representation channel wise which is then passed through two different batches. First branch is the Conv branch which contains depth wise and point wise convolution. Second branch is SSM branch which contains Parallel Efficient Vision Mamba (PEVM) block. Microscopic-Mamba has three variants in terms of parameters such as tiny, small, base. Tiny has 4.32M parameters, small has 4.97M parameters and base has 8.37M parameters.
From this section 3.5.2, we can see that MedMamba [21] was evaluated on various datasets using model variants like tiny, small, and base similar to Swin Transformer [130]. On the PAD-UFES-20 [157] dataset, Swin-T had a slightly better AUC of 0.830 compared to MedMamba-T’s 0.808, but it used more than twice the FLOPs (4.5G vs. 2.0G) and nearly double the parameters (27.5M vs. 14.5M). On the Cervical-US dataset, MedMamba-T reached a higher AUC of 0.952 compared to Swin-T’s 0.890, while staying more efficient. Similar results were seen on CPX-Ray [158], Kvasir [159] , and Fetal-Planes-DB [160], where MedMamba models gave better performance with fewer resources. while Swin Transformer sometimes achieves higher accuracy but at a higher computational cost.
3.5.3 Medical Image Restoration/ Reconstruction
@@
Restoration is an application in medical imaging which is used to improve the quality of images that may be corrupted or distorted due to factors such as noise, low resolution and blurring. Reconstruction is a mathematical process that converts raw medical data into target image. Table 5 provides an overview of Mamba based models applied in medical image restoration/reconstruction along with parameters, descriptions and the availability of code. Key Mamba-based architectures used in this application are explained as follows.
Table 5: Overview of Restoration Models
@@
MambaMIR [23] - Mamba MRI-CT image reconstruction
FDVM-Net [24] - Mamba Endoscopic image reconstruction
MambaDFuse [25] - Mamba Fused reconstruction with MRI,CT,PET images
FusionMamba [76] - DVSS Dynamic Feature Enhancement and Image fusion
@@
Zheng et al. [24] presents FDVM-Net which is designed for endoscopic exposure correction and it leverages frequency-domain reconstruction to achieve high-quality image restoration. To capture spatial features and global dependencies, FDVM-Net combines Mamba and convolutional blocks. This combination serves as the foundation for a dual-path network architecture, which separates processing of the image’s phase and amplitude information. Furthermore, FDVM-Net applies a frequency domain cross-attention module to improve the performance of network.
@@
Huang et al. [23] introduced MambaMIR for medical image reconstruction and uncertainty estimation. MambaMIR incorporates an Arbitrary-Masked State Space (AMSS) block with Monte Carlo dropout. AMSS is inspired by Mamba which comprises of AMS6 block, a gating linear layer, depth-wise convolution layer, and SiLU activation function. Residual connections used in MambaMIR ensures efficient and stable training. MambaDFuse introduced by Li et al. [25] incorporates Multimodal Mamba blocks which effectively merges features from different modality into a unified representation. This helps to obtain an informative fused image as output. MambaDFuse is a dual-phase model that integrates complementary information from diverse imaging modalities into a single image. MambaDFuse utilizes the combined strengths of convolutional layers and Mamba blocks to capture a spectrum of features. It ranges from low-level details to high-level semantic information.
@@
Xie et al. [76] proposed FusionMamba to address limitations of channel redundancy and limited local enhancement capabilities in existing image fusion methods. FusionMamba uses dynamic convolution and channel attention mechanisms to enhance the model’s ability to capture global context and local features within the images. FusionMamba leverages strengths of Mamba, dynamic feature enhancement and cross-modal fusion techniques. FusionMamba explores internal features and relationships between different image modalities. This results in dynamically enhanced texture details, better differences between modalities and improved ability to capture correlations while simultaneously reducing redundant information.
Section 3.5.3 presents the performance of the Mamba-based models for medical image restoration. To compare its effectiveness with transformer models such as SwinF [153], which uses Swin Transformer blocks for attention-guided cross-domain fusion. On the MRI-CT dataset, MambaDFuse achieved higher entropy (4.80 vs. 4.03) and slightly better structural similarity (SSIM: 0.67 vs. 0.65), while SwinF yielded a marginally higher visual information fidelity (VIF: 0.66 vs. 0.64). On the MRI-PET dataset, MambaDFuse again outperformed SwinF with higher entropy (4.91 vs. 3.90), VIF (0.65 vs. 0.63), and SSIM (0.52 vs. 0.51). For the MRI-SPECT dataset, both models achieved the same VIF (0.81), but MambaDFuse showed slight improvements in SSIM (0.64 vs. 0.63) and entropy (3.99 vs. 3.90). These results show that MambaDFuse performs slightly better than SwinF, offering better detail preservation and structure similarity across datasets while using Mamba’s lightweight architecture.
3.5.4 Medical Image Registration
@@
Image registration is the process involves aligning two or more images of the same scene captured at different times, from different viewpoints, or by different sensors. In medical image registration, this typically involves aligning a fixed volume, such as an intraoperative CT scan with a moving volume, such as a preoperative MRI to accurately overlay anatomical structures from different imaging modalities. This step is crucial for diagnosis, treatment planning, and monitoring in medical applications. Deformable registration refers to the alignment of images using transformation models that allow for local deformations.
Table 6: Overview of Registration Models
@@
MambaMorph [66] 7.59 Mamba 3D MRI-CT registration
VMambaMorph [75] 9.64 VMamba 3D MRI-CT registration
Deformable registration can handle complicated transformations which makes it appropriate for aligning MR and CT scans when the patient’s anatomy may have changed between scans. Table 6 describes registration models and their parameters along with information on their descriptions and code availability. Some mamba-based models relevant to this area include:
@@
Guo et al. [66] proposed a multi-modality deformable registration framework designed to process the medical MR-CT deformable registration dataset. Inspired by Chen et al. [161], which utilizes Swin Transformers in encoder of registration module, MambaMorph instead uses Mamba blocks to capture long-range spatial relationships while optimizing memory utilization. Before registration module, a simple fine-grained feature extractor (U-Net with one downsampling step) is utilized to maximize retention of local information.
@@
Wang et al. [75] transformed the 2D VSS of VMamba into a 3D volumetric feature processing framework. This hybrid architecture combined VMamba with CNN (U-Net), to accurately estimate the deformation field. To address the challenges posed by complex motion and diverse structures in image registration, VMambaMorph introduces a recursive registration framework. Furthermore, it employs a weight-sharing fine-grained feature extractor to extract features across divergent volumes. MambaMorph’s inadequacy in fully exploiting visual features from complex motion and diverse structures of images or volumes has been surpassed by VMambaMorph.
From this section 3.5.4, we can understand that VMambamorph is better than MambaMorph. While comparing VMambaMorph [75] with the transformer-based model TransMorph [161] on the SynthRAD dataset, VMambaMorph outperformed TransMorph in both Dice Score and Hausdorff Distance (HD). Specifically, TransMorph achieved a Dice score of 82.31% and a HD of 1.39, whereas VMambaMorph achieved a higher Dice score of 82.94% and a lower HD of 1.35, showing better overlap and reduced boundary error.
3.5.5 Miscellaneous